A) How much work must you do to push a 10kg block of steel across a steel table at a steady speed of 1.0 m/s for 3.0 s?

B) What's the power output while doing so?
1
Expert's answer
2020-05-04T12:39:37-0400

As per the given question,

mass of the object "(m)=10kg"

speed of the block "(v)=1.0 m\/sec"

time "(t)=3.0 sec"

a) work done = change in kinetic energy "(W)=\dfrac{mv^2}{2}=\dfrac{10\times 1}{2}=5J"

b) Power =rate of change work

"=\dfrac{5}{3} j/sec"